pandasurvive:

A flare was launched straight up from the ground with an initial velocity of 176 ft/s and returned to the ground after 11 s. The height of the flare t seconds after launch is modeled by the function f(t)=−16t2+176t . What is the maximum height of the flare, in feet?

callielovesyhuu:

One way to figure out the maximum height is to take the derivative of the equation. So, f'(t)=-32t+176 Find for what values of t f'(t) is 0, as that is the maximum/minimum of your function (when the slope of the tangent is 0). So 0=-32t+176 -32t=-176 t=5.5 s. So at 5.5 s is when a max height takes place. Plug it back into your original equation, and you should get: f(5.5)=-16(5.5)^2+176(5.5) f(5.5)=484 ft Another way to get it is by doing this: The maximum height must occur at it's vertex, which should be in the middle of your x-intercepts. Knowing that f(t)=-16t^2+176t, you can factor. 0=t(-16t+176) So the first x intercept is at 0, and the second is at x=11. To get the middle point, add up the x intercepts and divide it by 2, and you should get that x=5.5. Plug that back into your equation and you should have the maximum height.

Ultrilliam:

@pandasurvive

pandasurvive:

Just wondering if it was 16x^2 wouldnt it be 16x time 16x instead of 32x?

pandasurvive:

@callielovesyhuu

psirockin2:

f'(t)=-32t+176 Find for what values of t f'(t) is 0, as that is the maximum/minimum of your function (when the slope of the tangent is 0). So 0=-32t+176 -32t=-176 t=5.5 s. So at 5.5 s is when a max height takes place. Plug it back into your original equation, and you should get: f(5.5)=-16(5.5)^2+176(5.5) f(5.5)=484 ft I swear, teachers use this one for homework all the freaking time.

pandasurvive:

@psirockin2 just wondering are you sure?